OpenStudy (anonymous):

If f(x) = 4x - 6, what is f(-3)? 6 -49 -18 -6

OpenStudy (anonymous):

plug in -3 for x. f(-3)=4(-3)-6 f(-3)=-12-6 f(-3)=-18 so anwer C. can you answer mine?

OpenStudy (anonymous):

sure ill try

OpenStudy (anonymous):

thanks :) http://openstudy.com/study#/updates/4ff8b315e4b058f8b7632f31